# A Survey on Speech Large Language Models

### v2

@@ 8 Future Exploration
In terms of training strategies, innovation is also essential. Current training methods largely focus on pretraining and supervised fine-tuning, with reinforcement learning (RL) not yet being widely applied. However, when large language models are integrated into the model architecture, these training methods may not fully meet the requirements of training large models. Considering the diversity of inputs and outputs in large language models, exploring the incorporation of reinforcement learning strategies such as Proximal Policy Optimization (PPO) [54] can be beneficial. Introducing perturbations in supervised learning can help maintain the model’s multi-task performance. Additionally, to address hallucinations in large language models, techniques such as Reinforcement Learning with Human Feedback (RLHF) can be employed to apply ”soft” constraints [55], alleviating issues such as repetitive outputs and other hallucination phenomena.
@@ 9 Conclusion
This work presents a comprehensive exploration of Speech Large Language Models (Speech LLMs), focusing on their architecture, training strategies, and advancements in the field of Spoken Language Understanding (SLU). We examined the evolution from traditional methods to novel architectures that integrate Audio Feature Extraction, Multimodal Information Fusion, and LLM Inference, facilitating richer audio processing and deeper multimodal understanding. We summarized key methods for multimodal information fusion, including several approaches in audio-to-text modality conversion and combining audio and text feature spaces. Additionally, we discussed training strategies such as self-supervised learning and reinforcement learning that enable enhanced performance across various speech tasks. The survey highlights key improvements in Rich Audio Transcription and the potential for Cross-task Integration across SLU tasks. However, challenges such as the Dormancy of LLMs under certain conditions persist, underscoring the need for further innovation in training strategies and system design. By addressing these challenges and exploring future solutions, this work provides valuable insights into advancing Speech LLMs and their applications in multimodal contexts.

### v4

@@ 9 Future Exploration
In terms of training strategies, innovation is also essential. Current training methods largely focus on pretraining and supervised fine-tuning, with reinforcement learning (RL) not yet being widely applied. However, when large language models are integrated into the model architecture, these training methods may not fully meet the requirements of training large models. Considering the diversity of inputs and outputs in large language models, exploring the incorporation of reinforcement learning strategies such as Proximal Policy Optimization (PPO) [49] can be beneficial. Introducing perturbations in supervised learning can help maintain the model’s multitask performance. Additionally, to address hallucinations in large language models, techniques such as Reinforcement Learning with Human Feedback (RLHF) can be employed to apply ”soft” constraints [69], alleviating issues such as repetitive outputs and other hallucination phenomena.
@@ 10 Conclusion
This work presents a comprehensive exploration of Speech Large Language Models (Speech LLMs), focusing on their architecture, training strategies, and advancements in the field of Speech Understanding. We examined the evolution from traditional methods to novel architectures that integrate Audio Feature Extraction, Multimodal Information Fusion, and LLM Inference, facilitating richer audio processing and deeper multimodal understanding. We summarized key methods for multimodal information fusion, including several approaches in audio-to-text modality conversion and combining audio and text feature spaces. Additionally, we discussed training strategies such as self-supervised learning and reinforcement learning that enable enhanced performance across various speech tasks. The survey highlights key improvements in Rich Audio Transcription and the potential for Cross-task Integration across SLU tasks. However, challenges such as the Dormancy of LLM under certain conditions persist, underscoring the need for further innovation in training strategies and system design. By addressing these challenges and exploring future solutions, this work provides valuable insights into advancing Speech LLMs and their applications in multimodal contexts.